OpenStudy (ams98):

can someone help me!? How do I solve this? Algebra 2 ? medal and fan?

OpenStudy (anonymous):

probably

OpenStudy (ams98):

Simplify the given expression to radical form, justify each step by identifying the properties of rational exponents used. All work must be shown.\[x \frac{ 2 }{ 3}/ x \frac{ 4 }{ 9}\]

OpenStudy (ams98):

@DebbieG can you explain this too me?

OpenStudy (debbieg):

I don't see any exponents there. is it: \(\large \dfrac{x^{2/3}}{x^{4/9}}\) ??

OpenStudy (ams98):

yes i wasn't sure how to do that !

OpenStudy (debbieg):

Ok, use the rule that: \(\large \dfrac{x^{a}}{x^{b}}=x^{a-b}\)

OpenStudy (ams98):

i may be totally wrong but i got \[x ^{2/9}\]

OpenStudy (debbieg):

Yes, that's correct. Now you are supposed to put it into radical form. Remember, in a rational exponent, the numerator is the power, the denominator is the root.

OpenStudy (ams98):

So: \[\sqrt[9]{x^{2}}\]

OpenStudy (debbieg):

\[\Large x^{a/b}=\sqrt[b]{x^a}=\sqrt[b]{x}^a\] (It doesn't matter is the power goes inside, or outside, the radical... ) Yes, that's it!

OpenStudy (ams98):

okay thank you so much!!

OpenStudy (debbieg):

you're welcome. :)